OpenStudy (anonymous):

can someone tell me how to solve operations with radical expressions in the least complicated way?

OpenStudy (lgbasallote):

do you mean like \(\sqrt 2 + \sqrt 8 + \sqrt 18?\)

OpenStudy (anonymous):

example: 

OpenStudy (lgbasallote):

think of \(\sqrt 7\) as x so it's like 7x - 2x..do you know how to combine that?

OpenStudy (anonymous):

maybe haha my brain is fried. do you just subtract?

OpenStudy (lgbasallote):

yup! so....what's your answer?

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\(\Huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

Yay! Thanks!